# Static Charged Polytropic Spheres with a Cosmological Constant: Physical Acceptability and Trapped Orbits

### v1

@@ Abstract
We consider static charged fluid spheres with a cosmological constant. We assume a polytropic equation of state, $p\propto\rho^{\Gamma}$, and a power law charge distribution, $q\propto r^{n}$. Using this, we convert the generalised Tolman-Oppenheimer-Volkoff equation into a differential equation for the mass profile. By solving this equation numerically, we analyse both physical and geometric properties of these charged fluid spheres for different values of $n$ and $\Gamma$. By imposing subluminal sound speeds and energy conditions, we restrict ourselves to configurations that are physically acceptable. Then, within these physical models, we study internal trapping of circular geodesics and find the trapping regions in the $n$-$\Gamma$ parameter space. Going beyond the traditionally studied case of null geodesics, we consider orbits of charged and/or massive particles. We show that for neutral null particles (and only for them), their trapping depends purely on the properties of the space-time. In the other three cases, properties such as the particle’s own charge and/or energy also play a role. In general, we find that trapping of all types of particles is allowed for a broad range of $n$ and $\Gamma$.
@@ 1  Introduction
In the theory of general relativity (GR), static spherically symmetric fluid solutions have been historically significant due to their interesting mathematical properties as well as relevance in modelling compact objects. A common method to build these models is by taking an ansatz on the equation of state (EOS). Within this approach, a frequently used equation of states is the polytropic EOS [1, 2, 3, 4]. In this, the pressure ($p$) and density ($\rho$) are related through a simple power law relationship, $p\propto\rho^{\Gamma}$, where $\Gamma$ is known as the polytropic (or adiabatic) index. Such ‘polytropic solutions’ have also been generalised to include either a charge [5, 6, 7, 8, 9, 10] or a cosmological constant [11, 12, 13]. However, to the best of our knowledge, there have been no studies that involve both.
@@
Polytropic configurations have been used to describe neutron stars [14, 15, 16, 17], self bound stars (such as compact objects made of quark matter) [18, 19] and even non-compact objects such as dwarf galaxies and dark matter halos [20, 21]. In analyses of ultra-compact objects using exact solutions like the Tolman VII solution, it is common to study the phenomenon of the internal trapping of circular geodesics [22, 23, 24, 25, 26, 27]. The possibility of trapping of neutrinos or gravitational waves provides such studies with some astrophysical significance [25, 28]. Despite polytropes being used to model ultra-compact objects, trapping of geodesics in polytropic configurations has been studied in only a handful recent papers [29, 30, 31, 32].
@@
The goal of this paper is to extend the works mentioned above by providing an exposition to charged polytropic fluid spheres with a cosmological constant and the phenomenon of trapped geodesics within such configurations. To solve the polytropic system of equations, it is common to use the ansatz on the density profile due to Tooper [1] where, $\rho\propto\theta^{\frac{1}{1-\Gamma}}$. This leads to a relativistic analogue of the Lane-Emden equation [1, 11, 29], which can then be solved. Here, we take a different route and reformulate the generalised Tolman-Oppenheimer-Volkoff (TOV) equation in terms of the mass profile as done previously in [33]. We extend the treatment in [33] to include charge and a cosmological constant. This leads to a highly non-linear differential equation with two unknowns – the mass profile and the charge distribution. We assume a power law charge distribution, $q\propto r^{n}$, which has been extensively used in the literature [34, 35, 36, 37, 38]. Then, for a range of values of $\Gamma$ and $n$, we solve this differential equation numerically to find the mass profile. Using the mass profile, we calculate the physical properties of the fluid such as density, pressure and sound speed. We also analyse energy conditions and filter out the physically acceptable configurations. After the physical properties, we calculate the geometric properties such as metric functions and spatial curvature for these cases.
@@
We then turn our attention to the phenomenon of internal trapping of circular geodesics. For a range of values of the central density, $\rho_{0}$, central pressure, $p_{0}$, and total charge, $Q$, we explore the $n$-$\Gamma$ parameter space to look for values that allow such trapping. We do this for all four possible types of particles: neutral null, neutral timelike, charged null, and charged timelike. Only for the case of neutral null particles (such as photons and gravitons), whether a circular geodesic will be trapped is determined entirely by the gravitational potential. For the other three types, properties such as charge and energy of the particles also play a role. We analyse their effect on the trapping as well. In general, we find that trapping occurs for all the cases for a range of central density and total charge values that correspond to ultra-compact objects as well as a range of $\Gamma$ and $n$. We also find that the bounds on the $n$-$\Gamma$ parameter space for different central densities do not differ in a significant manner but increasing the total charge and the cosmological constant is detrimental to having trapped orbits.
@@ 2.1  Solutions to the Field Equations
Before we present our polytropic models, we would like to briefly discuss their place in the landscape of static spherically symmetric charged perfect fluid solutions of the EM-$\Lambda$ field equations. Since the system of equations is not closed, additional assumptions are necessary. A classification scheme was presented in [40], based on which variables are specified to close the system. However, we recognise that such assumptions are usually also based on a physical modelling methodology rather than only the purely mathematical issue of closure. For example, we choose to work with a polytropic EOS and a power law charge distribution. Considering this, we identify three broad classes of such modelling methodologies: (I) those that make assumptions on the matter variables (density, pressure, charge, or an EOS); (II) those that make assumptions on the geometric data (one or both metric functions, a relation between metric functions, an embedding or symmetry condition, or a generating function); and (III) those that make assumptions on both the geometry and the matter sector simultaneously. Further, sub-classes can be defined according to the specific assumptions that go in construction of the solutions.
@@ 2.2  Charged Polytropic Solutions
We consider the range, $1\leq\Gamma\leq 5$, since the polytropic index for compact objects falls in this range [42, 43]. We note that the above equation becomes singular at centre for $n<{\frac{1}{2}}$ and everywhere for $n={\frac{1}{2}}$. To avoid this, we choose the range, $1\leq n\leq 5$. We also vary the values of $\rho_{0}$ and $Q$ and present several cases. The values of these two parameters are taken to be their typical values for ultra-compact objects, which in geometric units become, $\rho_{0}\sim 10^{-9}$ m-2 and $Q\sim 10^{3}$ m [5, 6] while $\kappa$ is calculated by fixing the ratio, $\frac{p_{0}}{\rho_{0}}=0.2$. We use this value because it satisfies the constraint, $\frac{p_{0}}{\rho_{0}}<\frac{1}{\Gamma}$ that comes from causality [11, 30]. Further, the cosmological constant is fixed by cosmology [44] while we take $r_{b}=10^{4}$ m which is a typical size of a neutron star. We plot the mass profile for different values of these parameter in figure 1 below.
We find that for these ranges of different parameters, the mass distribution is relatively well-behaved. Corresponding to different values of $Q$, the generalised Buchdahl limit [45] on the total inertial mass can be calculated to be between $4\times 10^{3}$ m to $6\times 10^{3}$ m. The total mass, $m(r_{b})$, falls in this range for all the cases. For $\rho_{0}\gtrsim\times 10^{-8}$ and $Q\gtrsim 10^{4}$, distribution slowly ceases behaving well as seen in the bottom row of the plot. Further, in all of the well-behaved cases, the models obey the generalised Buchdahl limit [45] for $Q\sim 10^{3}$. We will show in the proceeding sections that these cases also follow the standard physical acceptability criteria.
Figure 1: The mass profile, $m(r)$, of various charged polytropic configurations (different values of $n$ and $\Gamma$) obtained by solving equation (2.20). The central density, $\rho_{0}$, and total charge, $Q$, correspond to their typical values for ultra compact objects [46, 5, 6]. The boundary radius is, $r_{b}=10^{4}$ m and $\Lambda=10^{-52}$ m-2.
@@ 3.1  Density and Pressure
The pressure can be found by using the density profile above in equation (2.15). We plot the density and pressure below.
@@
Figures 2 and 3 show the radial profiles of the energy density and pressure of the charged polytropic fluid for different values of $\Gamma$ and $n$. For most of the values where $1<n,\Gamma<5$, both the density and pressure are monotonically decreasing and the pressure tends to vanish at the boundary. However, not all configurations exhibit these features. We filter such cases out, along with others that do not follow additional physical acceptability criteria in the following sections.
@@ 3.2  Sound Speed and Bulk Modulus
We plot the sound speed and the bulk modulus in figures 4 and 5 below. We find that causality id respected in most of the cases as well as the sound speed is monotonically decreasing.
@@ 3.3  Energy Conditions
We plot these energy conditions one by one below in figures 6, 7, 8, and 9, respectively.
@@
Figure 10: The regions in the $n$-$\Gamma$ parameter space for which the corresponding charged polytropic static spherically symmetric fluid configurations follow the basic physically acceptability criteria.
@@ 4.1  Metric Functions
Figure 11: The radial profile of the metric function, $\textrm{e}^{2\Phi}$.
Figure 12: The radial profile of the metric function, $\textrm{e}^{2\Psi}$.
@@ 4.2  Spatial Curvature
The function, $k(r)$, above is directly proportional to the scalar curvature of the $t={\rm const.}$ hypersurfaces. In other words, it measures the spatial curvature of the configuration. In figure 13, we plot the spatial curvature as a function of the radial coordinate, $r$. We find that for our chosen values of $\rho_{0}$ and $Q$, the spatial geometry remains spherical (i.e., $k(r)>0$) throughout the extent of the fluid sphere for $1<n,\Gamma<5$. However, in the case of a negative cosmological constant, the spatial curvature may take negative values (only when $|\Lambda|\sim\rho_{0}$).
Figure 13: The radial profile of the spatial curvature, $k(r)$, for various polytropic configurations. The spatial curvature remains positive, i.e., the spatial geometry remains spherical for all of our models ($1<n,\Gamma<5$) constructed with realistic values of $\rho_{0}$ and $Q$.
@@ 5  Effective Potential and Trapped Circular Geodesics
where, $\mu=0\ {\rm and}\ -1$ for null and timelike particles, respectively.
@@ 5.1  Trapping Regions
With the expression for the effective potential and trapping conditions derived in the previous section, we now have everything necessary to check the existence of trapped orbits. We do this for all the four different particle types listed in table 1. We explore the $n$-$\Gamma$ parameter space and find regions that allow for trapped orbits. We refer to them as ‘trapping regions’. Below, we present the plots for the effective potential and the trapping regions in all the four cases. All the plots are made with $\Lambda=10^{-52}$ m-2. To analyse the effect of changing $\Lambda,\ Q,\ r_{b},\ \mathcal{E}$, and $\frac{e}{\mathcal{E}}$ on the trapping regions, we also generate animations sweeping through various values values of these parameters [^3] These animations could be found at this link: https://github.com/astornelli/CPS..
We find that trapping regions exist for all types of particles for typical values of $\rho_{0}$, $\Lambda$, $Q$, and $r_{b}$. For neutral null particles, trapping regions become smaller as $Q$ or $\Lambda$ increase. For charged and/or massive particles, the energy per unit mass, $\mathcal{E}$ and the charge to energy ratio, $\frac{e}{\mathcal{E}}$ affects the trapping regions as well. We find that trapping regions do not exist for neutral timelike particles for $\mathcal{E}<1$ while increasing $\mathcal{E}$ beyond this does not have a significant effect. For charged timelike particles, increasing $\frac{e}{\mathcal{E}}$ makes the trapping regions smaller. For all the four cases, the trapping regions become bigger with increasing $r_{b}$.
@@
Figure 15: The trapping regions for neutral, null particles for various values of $\rho_{0}$ and $Q$.
@@
Figure 17: The trapping regions for neutral, timelike particles for various values of $\rho_{0}$ and $Q$.
@@
Figure 19: The trapping regions for charged, null particles for various values of $\rho_{0}$ and $Q$. We have used, $\frac{e}{\mathcal{E}}=6$.
@@
Figure 21: The trapping regions for charged, timelike particles for various values of $\rho_{0}$ and $Q$. We have used, $\frac{1}{\mathcal{E}^{2}}=0.04$ and $\frac{e}{\mathcal{E}}=8.3$.
@@ 6  Summary and Discussion
Then, we turned our attention to the phenomenon of internal trapping of circular geodesics which has been previously studied mainly for neutral, massless particles in uncharged polytropic configurations [30]. Here, we included the massive and charged cases as well as included a charge into the fluid. Using the effective potential for these cases, we showed that trapping is allowed for all types of particles for a broad range of $n$ and $\Gamma$. We found that higher values of $Q$ reduced the possibility of trapping while higher values of $r_{b}$ increased it. On the other hand, the cosmological constant enters the geodesic equations only via the metric function, $\textrm{e}^{2\Phi}$. For the currently accepted astrophysical value of $10^{-52}$ m-2, it has negligible effect on geodesics. As was the case with physical acceptability, for $\Lambda$ to become relevant for geodesic trapping, it must be $\sim 10^{-9}$, which is comparable to $\rho_{0}$. In such cases, the cosmological constant leads to even smaller trapping regions.
@@
We noted that, unlike in the case of neutral, massless particles, the effective potential for massive/charged particles also depends on properties such as specific energy, $\mathcal{E}$, and specific charge, $e$. For the case of neutral, massive particles, we found that trapped geodesics exist only for $\mathcal{E}>1$. Increasing $\mathcal{E}$ beyond this does not have any qualitative effect on the trapping regions. This makes sense since from table 1, we can see that $\mathcal{E}$ must be greater than unity for the trapping condition to be satisfied since other terms are smaller than unity. For charged, massless particles, trapping takes place only for $\frac{e}{\mathcal{E}}\lesssim 10$ while decreasing the value of this ratio leads to larger trapping regions. For charged, massive particles, both $\mathcal{E}$ and $e$ play a role where their effect on the trapped regions remains qualitatively the same as in the other mentioned above.
@@
It is important to note here that the actual physical trajectories of the test particles will also depend on additional physical effects such as radiation backreaction (for charged particles) and scattering. Therefore, our results here should be considered only a first step towards a more comprehensive analysis of trapped orbits in charged polytropic configurations. Moreover, charged, massless particles have not been observed in nature but do exist in various contexts in the literature [48, 49, 50, 51, 52, 53, 54]. Nevertheless, the possibility of trapped orbits means that this work might prove relevant for the study of neutrino trapping [24, 55, 25] and black hole mimickers [56, 57, 30].
@@
Finally, we remark on other possible constructions of static charged fluid spheres with $\Lambda$. One could construct models where, instead of a power law charge distribution, one assumes that the charge density is proportional to the matter density as done in [5, 7, 8]. This would fall into class Ic under the scheme described in section 2.1. Other classes of solutions could also be derived. A generalised relativistic Lane-Emden type system with charge and $\Lambda$ can also be derived in the same manner as [58, 11]. In addition to these, a direct follow up to the current work may include stability analysis and anisotropic extensions of our models [59, 60, 61]. We keep these issues, particularly the last two, for future work.

### v2

@@ Abstract
We consider static charged fluid spheres with a cosmological constant. We assume a polytropic equation of state, $p\propto\rho^{\Gamma}$, and a power law charge distribution, $q\propto r^{n}$. Using this, we convert the generalised Tolman-Oppenheimer-Volkoff equation into a differential equation for the mass profile. By solving this equation numerically, we analyse both physical and geometric properties of charged polytropic fluid spheres for different values of $n$ and $\Gamma$. By imposing subluminal sound speeds and energy conditions, we restrict ourselves to configurations that are physically acceptable. Then, within these physical models, we study internal trapping of circular geodesics and find the trapping regions in the $n$-$\Gamma$ parameter space. Going beyond the traditionally studied case of null geodesics, we consider orbits of charged and/or massive particles as well. We show that for neutral null particles (and only for them), the possibility of internal trapping is determined purely through geometry. In the other three cases, properties such as the particle’s own charge and/or energy also play a role. In general, we find that trapping of all types of particles is allowed for a broad range of $n$ and $\Gamma$.
@@ 1  Introduction
In the theory of general relativity (GR), static spherically symmetric fluid solutions have been historically significant due to their interesting mathematical properties as well as relevance in modelling compact objects. A common method to build these models is by taking an ansatz on the equation of state (EOS). Within this approach, a frequently used equation of state is the polytropic EOS [1, 2, 3, 4]. In this, the pressure ($p$) and density ($\rho$) are related through a simple power law relationship, $p\propto\rho^{\Gamma}$, where $\Gamma$ is known as the polytropic (or adiabatic) index. Such ‘polytropic solutions’ have also been generalised to include either a charge [5, 6, 7, 8, 9, 10] or a cosmological constant [11, 12, 13]. However, to the best of our knowledge, there have been no studies that include both.
@@
Polytropic configurations have been used to describe neutron stars [14, 15, 16, 17], self bound stars (such as compact objects made of quark matter) [18, 19] and even non-compact objects such as dwarf galaxies and dark matter halos [20, 21]. In analyses of ultra-compact objects using exact solutions like the Tolman VII solution, it is common to study the phenomenon of the internal trapping of circular geodesics [22, 23, 24, 25, 26, 27]. The possibility of trapping of neutrinos or gravitational waves provides such studies with some astrophysical significance [25, 28]. Despite polytropes being used to model ultra-compact objects, trapping of geodesics in polytropic configurations has been studied in only a handful of recent papers [29, 30, 31, 32].
@@
The goal of this paper is to extend the works mentioned above by providing an exposition to charged polytropic fluid spheres with a cosmological constant and the phenomenon of trapped circular geodesics within such configurations. To solve the polytropic system of equations, it is common to use the ansatz on the density profile due to Tooper [1] where, $\rho\propto\theta^{\frac{1}{1-\Gamma}}$. This leads to a relativistic analogue of the Lane-Emden equation [1, 11, 29], which can then be solved. Here, we take a different route and reformulate the generalised Tolman-Oppenheimer-Volkoff (TOV) equation in terms of the mass profile as done previously in [33]. We extend the treatment in [33] to include charge and a cosmological constant. This leads to a highly non-linear differential equation with two unknowns – the mass profile and the charge distribution. We assume a power law charge distribution, $q\propto r^{n}$, which has been extensively used in the literature [34, 35, 36, 37, 38]. Then, for a range of values of $\Gamma$ and $n$, we solve this differential equation numerically to find the mass profile. Using the mass profile, we calculate the physical properties of the fluid such as density, pressure and sound speed. We also analyse energy conditions and filter out the physically acceptable configurations. After the physical properties, we calculate the geometric properties such as metric functions and spatial curvature for these cases.
@@
We then turn our attention to the phenomenon of internal trapping of circular geodesics. For a range of values of the central density, $\rho_{0}$, central pressure, $p_{0}$, and total charge, $Q$, we explore the $n$-$\Gamma$ parameter space to look for values that allow trapped orbits. We do this for all four possible types of particles: neutral null, neutral timelike, charged null, and charged timelike. Only for the case of neutral null particles (such as photons and gravitons), whether a circular geodesic will be trapped is determined entirely by the gravitational potential. For the other three types, properties such as charge and energy of the particles also play a role. We analyse the effect of these properties on trapping as well. In general, we find that trapping occurs for all the cases for a range of central density and total charge values that correspond to ultra-compact objects as well as a range of $\Gamma$ and $n$. We also find that the bounds on the $n$-$\Gamma$ parameter space for different central densities do not differ in a significant manner but increasing the total charge and the cosmological constant is detrimental to having trapped orbits.
@@ 2.1  Solutions to the Field Equations
Before we present our polytropic models, we will briefly discuss their place in the landscape of static spherically symmetric charged perfect fluid solutions of the EM-$\Lambda$ field equations. Since the system of equations is not closed, additional assumptions are necessary. A classification scheme was presented in [40], based on which variables are specified to close the system. However, we recognise that such assumptions are usually also based on a physical modelling methodology rather than only the purely mathematical issue of closure. For example, we choose to work with a polytropic EOS and a power law charge distribution. Considering this, we identify three broad classes of such modelling methodologies: (I) those that make assumptions on the matter variables (density, pressure, charge, or an EOS); (II) those that make assumptions on the geometric data (one or both metric functions, a relation between metric functions, an embedding or symmetry condition, or a generating function); and (III) those that make assumptions on both the geometry and the matter sector simultaneously. Further, sub-classes can be defined according to the specific assumptions that go in construction of the solutions.
@@ 2.2  Charged Polytropic Solutions
We consider the range, $1\leq\Gamma\leq 5$, since the polytropic index for compact objects falls in this range [42, 43]. We note that the above equation becomes singular at centre for $n<{\frac{1}{2}}$ and everywhere for $n={\frac{1}{2}}$. To avoid this, we choose the range, $1\leq n\leq 5$. We also vary the values of $\rho_{0}$ and $Q$ and present several cases. The values of these two parameters are taken to be their typical values for ultra-compact objects, which in geometric units become, $\rho_{0}\sim 10^{-9}$ m-2 and $Q\sim 10^{3}$ m [5, 6] while $\kappa$ is calculated by fixing the ratio, $\frac{p_{0}}{\rho_{0}}=0.2$. We use this value because it satisfies the constraint, $\frac{p_{0}}{\rho_{0}}<\frac{1}{\Gamma}$ that comes from causality [11, 30]. Further, the cosmological constant is fixed by cosmology [44] while we take $r_{b}=10^{4}$ m which is a typical size of a neutron star. We plot the mass profile for different values of these parameter in figure 1 below. We find that for these ranges of different parameters, the mass distribution is relatively well-behaved except for $\rho_{0}\gtrsim 10^{-8}$ and $Q\gtrsim 10^{4}$ as can be seen in the bottom row of the plot. Corresponding to different values of $Q$, the generalised Buchdahl limit [45] on the total inertial mass can be calculated to be between $4\times 10^{3}$ m to $6\times 10^{3}$ m. The total mass, $m(r_{b})$, falls in this range for all the well-behaved cases. We will show in the proceeding sections that these cases also follow the standard physical acceptability criteria.
Figure 1: The mass profile, $m(r)$, of various charged polytropic configurations (different values of $n$ and $\Gamma$) obtained by solving equation (2.20). The central density, $\rho_{0}$, and total charge, $Q$, correspond to their typical values for ultra-compact objects [46, 5, 6]. The boundary radius is, $r_{b}=10^{4}$ m and $\Lambda=10^{-52}$ m-2.
@@ 3.1  Density and Pressure
The pressure can be found by using the density profile above in equation (2.15). We plot the radial profile of the energy density and pressure for different values of $\Gamma$ and $n$ in figures 2 and 3 below. For most of the values where $1\leq n,\Gamma\leq 5$, both the density and pressure are monotonically decreasing and the pressure tends to vanish at the boundary. However, not all configurations exhibit these features. We filter such cases out, along with others that do not follow additional physical acceptability criteria in the following sections.
@@ 3.2  Sound Speed and Bulk Modulus
We plot the sound speed and the bulk modulus in figures 4 and 5 below. We find that causality is respected in most of the cases as well as the sound speed is monotonically decreasing.
@@ 3.3  Energy Conditions
We plot these energy conditions in figures 6, 7, 8, and 9, respectively.
@@
Figure 10: The physically acceptable regions in the $n$-$\Gamma$ parameter space for a given value of central density, $\rho_{0}$ and total charge, $Q$.
@@ 4.1  Metric Functions
Figure 11: The radial profile of the metric function, $\textrm{e}^{2\Phi(r)}$. The function remains smooth for most typical values of $\rho_{0}$ and $Q$. From the bottom row, one can see that for large values ($\rho_{0}>10^{-8}$ and $Q>10^{4}$), discontinuities may be present in the metric function.
Figure 12: The radial profile of the metric function, $\textrm{e}^{2\Psi(r)}$. The function remains smooth for most typical values of $\rho_{0}$ and $Q$.
@@ 4.2  Spatial Curvature
The function, $k(r)$, above is directly proportional to the scalar curvature of the spatial ($t={\rm const.}$) hypersurfaces. In figure 13, we plot the spatial curvature as a function of the radial coordinate, $r$. We find that for our chosen values of $\rho_{0}$ and $Q$, the spatial geometry remains spherical (i.e., $k(r)>0$) throughout the extent of the fluid sphere for $1\leq n,\Gamma\leq 5$. However, in the case of a negative cosmological constant, the spatial curvature may take negative values but only when its value is comparable to the central density ($|\Lambda|\sim\rho_{0}$).
Figure 13: The radial profile of the spatial curvature, $k(r)$, for various polytropic configurations. The spatial curvature remains positive, i.e., the spatial geometry remains spherical for all of our models ($1\leq n,\Gamma\leq 5$) constructed with realistic values of $\rho_{0}$ and $Q$.
@@ 5  Effective Potential and Trapped Circular Geodesics
where, $\mu=0\ {\rm and}\ -1$ correspond to null and timelike particles, respectively.
@@ 5.1  Trapping Regions
With the expression for the effective potential and trapping conditions derived in the previous section, we now have everything necessary to check the existence of trapped orbits. We do this for all four types of particle listed in table 1. We explore the $n$-$\Gamma$ parameter space and find regions that allow for trapped orbits, which we refer to as, ‘trapping regions’. Below, we present the plots for the effective potential and the trapping regions in all the four cases. All the plots are made with $\Lambda=10^{-52}$ m-2. To analyse the effect of changing $\Lambda,\ Q,\ r_{b},\ \mathcal{E}$, and $\frac{e}{\mathcal{E}}$ on the trapping regions, we also generate animations sweeping through various values values of these parameters [^3] These animations could be found at this link: https://github.com/astornelli/CPS.. The plots below can be considered as one of the frames in the corresponding animations corresponding to a fixed value of a given parameter that is being swept though.
We find that trapping regions exist for all types of particles for typical values of $\rho_{0}$, $\Lambda$, $Q$, and $r_{b}$. For neutral null particles, trapping regions become smaller as $Q$ or $\Lambda$ increase. For charged and/or massive particles, the energy per unit mass, $\mathcal{E}$ and the charge to energy ratio, $\frac{e}{\mathcal{E}}$ affects the trapping regions as well. We find that trapping regions do not exist for neutral timelike particles for $\mathcal{E}<1$ while increasing $\mathcal{E}$ beyond this does not have a significant effect. For charged timelike particles, increasing $\frac{e}{\mathcal{E}}$ makes the trapping regions smaller. For all four cases, the trapping regions become larger with increasing $r_{b}$.
@@
Figure 15: The trapping regions for neutral, null particles for various values of $\rho_{0}$ and $Q$. The region shaded blue corresponds to existence of only stable orbits while purple corresponds to both stable and unstable orbits.
@@
Figure 17: The trapping regions for neutral, timelike particles for various values of $\rho_{0}$ and $Q$. For the chosen range of $n,\ \Gamma,\ \rho_{0},\ Q$, only stable orbits exist.
@@
Figure 19: The trapping regions for charged, null particles for various values of $\rho_{0}$ and $Q$. We have used, $\frac{e}{\mathcal{E}}=6$. The region shaded blue corresponds to existence of only stable orbits while purple corresponds to both stable and unstable orbits.
@@
Figure 21: The trapping regions for charged, timelike particles for various values of $\rho_{0}$ and $Q$. We have used, $\frac{1}{\mathcal{E}^{2}}=0.04$ and $\frac{e}{\mathcal{E}}=8.3$. The region shaded blue corresponds to existence of only stable orbits while purple corresponds to both stable and unstable orbits.
@@ 6  Summary and Discussion
Then, we turned our attention to the phenomenon of internal trapping of circular geodesics which has been previously studied mainly for neutral, massless particles in uncharged polytropic configurations [30]. Here, we included the massive and charged cases as well as included a charge into the fluid. Using the effective potential for these cases, we showed that trapping is allowed for all types of particles for a broad range of $n$ and $\Gamma$. We found that higher values of $Q$ reduced the possibility of trapping while higher values of $r_{b}$ increased it. On the other hand, the cosmological constant enters the geodesic equations only via the metric function, $\textrm{e}^{2\Phi}$. For the currently accepted astrophysical value of $10^{-52}$ m-2, it has negligible effect on geodesics. As was the case with physical acceptability, for $\Lambda$ to become relevant for geodesic trapping, it must be comparable to $\rho_{0}$. In such cases, the cosmological constant leads to even smaller trapping regions.
@@
We noted that, unlike in the case of neutral, massless particles, the effective potential for massive/charged particles also depends on properties such as specific energy, $\mathcal{E}$, and specific charge, $e$. For the case of neutral, massive particles, we found that trapped geodesics exist only for $\mathcal{E}>1$. Increasing $\mathcal{E}$ beyond this does not have any qualitative effect on the trapping regions. This makes sense since from table 1, we can see that $\mathcal{E}$ must be greater than unity for the trapping condition to be satisfied since other terms are smaller than unity. For charged, massless particles, trapping takes place only for $\frac{e}{\mathcal{E}}\lesssim 10$ while decreasing the value of this ratio leads to larger trapping regions. For charged, massive particles, both $\mathcal{E}$ and $e$ play a role where their effect on the trapped regions remains qualitatively the same as in the other cases mentioned above.
@@
It is important to note here that in reality, trajectories of the test particles will also depend on additional physical effects such as radiation backreaction (for charged particles) and scattering. Therefore, our results here should be considered only a first step towards a more comprehensive analysis of trapped orbits in charged polytropic configurations. Moreover, charged, massless particles have not been observed in nature but do exist in various contexts in the literature [48, 49, 50, 51, 52, 53, 54]. Nevertheless, the possibility of trapped orbits means that this work might prove relevant for the study of neutrino trapping [24, 55, 25] and black hole mimickers [56, 57, 30].
@@
Finally, we remark on other possible constructions of static charged fluid spheres with $\Lambda$. One could construct models where, instead of a power law charge distribution, one assumes that the charge density is proportional to the matter density as done in [5, 7, 8]. This would fall into class Ic under the scheme described in section 2.1. Other classes of solutions could also be derived. A generalised relativistic Lane-Emden type system with charge and $\Lambda$ can also be derived in a similar manner as [58, 11]. In addition to these, a direct follow up to the current work may include stability analysis and anisotropic extensions of our models [59, 60, 61]. We keep these issues, particularly the last two, for future work.